# Transforming Image Super-Resolution: A ConvFormer-based Efficient Approach

### v1

@@ III Proposed Method
III-A Network Architecture
In Figure 2, we illustrate the proposed CFSR framework, which encompasses three pivotal stages: shallow feature extraction, deep feature extraction, and the image reconstruction module. The shallow feature extraction phase is designed to distill low-level image features, such as edges, textures, and fine-grained details, from the input image and map them into a latent space. These features are important for preserving the local structure and details of the image. In the deep feature extraction stage, the model extracts higher-level, more compact feature representations. It captures texture and structure information, which is essential for recovering lost details and enhancing the clarity of the image. Utilizing both shallow and deep features, the image reconstruction module is capable of generating high-resolution images. In the following, we will present the details of these three components.
@@ III-A Network Architecture
Deep feature extraction. Then we use a stack of two basic residual blocks (BRB), which contains several ConvFormer layers (CFL), and a $3\times 3$ convolution layer is added at the end of the BRB to aggregate the local features. Specifically, the detailed implementation of BRB and CFL are presented in Figure 2(a) and Figure 2(b), respectively. This process of deep feature extraction is formulated as:
@@ III-B ConvFormer Layer
Large kernel mixer. Self-attention is a powerful feature extractor, but its high computational cost makes it impractical for real-time application. Recent studies have demonstrated the effectiveness of employing large kernel convolutions[31]. In Table I, a detailed comparison of the computational complexity between multi-head self-attention (MHSA) [22], local window self-attention (LWSA) [23], and large kernel convolution (LK) [31] is exhibited.
Here the $H,W,C$, and $K$ represent the height, width, channel dimension, and kernel (local window) size, respectively. The comparison reveals that, given identical window and kernel sizes, the computational complexity of the LK token mixer is significantly less than that of MHSA. The complexity of LK approximates to half that of LWSA, while simultaneously exhibiting a reduction in parameter count, where $K<<C$. Consequently, large kernel convolutions offer a more resource-efficient choice for the design of lightweight models. In this paper, we propose a simple yet effective feature mixer module as presented in Figure 2(c). Here we take a $1\times 1$ convolution followed by a large kernel convolution as the feature mixing gate. Feature extraction of our LK mixer is formulated as follows:
@@
$F_{gate}$ $=\text{Conv}_{1\times 1}(\text{DwConv}_{k\times k}(F)),$
$F_{out}$ $=V\odot F_{gate},$
@@
Edge-preserving feed-forward network. In the feed-forward network (FFN) in a vanilla Transformer unit, previous studies have enhanced it by integrating depth-wise convolution, thereby improving the local feature ensemble[34, 33]. Considering that SISR, being an ill-posed problem, aims at learning an inversion from LR to HR, where the high-frequency information is crucial to it [53]. To obtain more high-frequency information in latent features, we propose an edge-preserving feed-forward network (EFN), by our Edge-preserving Depth-Wise Convolution (EDC), as illustrated in Figure 2(d). This allows a best of both words for a richer high-frequency information while maintaining local feature ensemble. The implementation of the proposed EFN is as follows:
@@
$F_{2}$ $=\text{GELU}(F_{1}),$ (8)
$F_{EDC}$ $=\text{EDC}(F_{2}),$ (9)
$F_{3}$ $=\text{ReLU}(F_{EDC}),$ (10)
$F_{4}$ $=\text{Conv}_{1\times 1}(F_{EDC}),$ (11)
@@
Detailed implementation of our EDC is presented in Figure 3. It takes a multi-branch structure, containing a standard depth-wise convolution (DwConv) and three DwConvs with pre-defined gradient kernels. Denote $K_{3\times 3}\in\mathbb{R}^{C\times 1\times 3\times 3}$ and $B_{3\times 3}$ the learnable kernel weights and bias for the vanilla DwConv, where $C$ presents the output channels and 3 is the spatial size. The feature extraction is formulated as:
@@
where *** presents the depth-wise convolution operation.
@@
$K_{Lap4}=\left[\begin{array}[]{ccc}0&+1&0\\ +1&-4&+1\\ 0&+1&0\par\end{array}\right],~{}K_{Lap8}=\left[\begin{array}[]{ccc}+1&+1&+1\\ +1&-8&+1\\ +1&+1&+1\end{array}\right],$ (15)
@@
TABLE II: Quantitative comparison with some state-of-the-art SR approaches on five widely used benchmark datasets. Mult-Adds is evaluated on a $1280\times 720$ HR image. Results of ours are in bold. One can find that our CFSR achieves SOTA performance among existing advanced CNN-based methods and significantly bridge the gap between Transformer-based methods.
@@
SRCNN [4] 8K 52.7G 36.66/0.9067 32.45/0.9067 31.36/0.8879 29.50/0.8946 35.60/0.9663
VDSR [5]  [^6] K 612.6G 37.53/0.9587 33.03/0.9124 31.90/0.8960 30.76/0.9140 37.22/0.9750
LapSRN [55] 251K 29.9G 37.52/0.9591 32.99/0.9124 31.80/0.8952 30.41/0.9103 37.27/0.9740
IDN [41] 553K 124.6G 37.83/0.9600 33.30/0.9148 32.08/0.8985 31.27/0.9196 38.01/0.9749
CARN[14] 1592K 222.8G 37.76/0.9590 33.52/0.9166 32.09/0.8978 31.92/0.9256 38.36/0.9765
SRResNet[56] 1,370K 341.7G 38.05/0.9607 33.64/0.9178 32.22/0.9002 32.23/0.9295 38.05/0.9607
sLWSR [57] 322K 20.6G 31.87/0.8910 28.06/0.7740 27.46/0.7320 25.65/0.772 -/-
@@
DDistill-SR[59] 414K 128.0G 38.03/0.9606 33.61/0.9182 32.19/0.9000 32.18/0.9286 38.94/0.9777
RFDN[60] 534K 95.0G 38.05/0.9606 33.68/0.9184 32.16/0.8994 32.12/0.9278 38.88/0.9773
@@
ELAN-light[61] 582K 168.4G 38.17/0.9611 33.94/0.9207 32.30/0.9012 32.76/0.9340 39.11/0.9782
VDSR [5]  [^6] K 612.6G 33.66/0.9213 29.77/0.8314 28.82/0.7976 27.14/0.8279 32.01/0.9340
LapSRN [55] 502K 90.0G 33.81/0.9220 29.79/0.8325 28.82/0.7980 27.07/0.8275 32.21/0.9350
CARN[14] 1,592K 118.8G 34.29/0.9255 30.29/0.8407 29.06/0.8034 28.06/0.8493 33.50/0.9440
SRResNet[56] 1,554K 190.2G 34.41/0.9274 30.36/0.8427 29.11/0.8055 28.20/0.8535 33.54/0.9448
@@
ELAN-light [61] 590K 75.7G 34.61/0.9288 30.55/0.8463 29.21/0.8081 28.69/0.8624 34.00/0.9478
SRCNN[4] 8K 52.7G 30.48/0.8626 27.50/0.7513 26.90/0.7101 24.52/0.7221 27.58/0.8555
VDSR[5]  [^6] K 612.6G 31.35/0.8838 28.01/0.7674 27.29/0.7251 25.18/0.7524 28.83/0.8870
LapSRN [55] 813K 149.4G 31.54/0.8852 28.09/0.7700 27.32/0.7275 25.21/0.7562 29.09/0.8900
IDN[41] 553K 32.3G 31.82/0.8903 28.25/0.7730 27.41/0.7297 25.41/0.7632 29.41/0.8942
CARN [14] 1592K 90.9G 32.13/0.8937 28.60/0.7806 27.58/0.7349 26.07/0.7837 30.47/0.9084
SRResNet[56] 1,518K 146G 32.17/0.8951 28.61/0.7823 27.59/0.7365 26.12/0.7871 30.48/0.9087
@@ IV-A Experimental Setup
Comparison Methods We compare the proposed CFSR with state-of-the-art efficient SR approaches, including SRCNN [4], VDSR [5], LapSRN [55], IDN [41], CARN [14], SRResNet[56], sLWSR[57], IMDN [15], LatticeNet[62], LAPAR[16], SMSR[17], ECBSR[19], DRSAN[58], PAN [18], DDistill[59], RFDN [60], ShuffleMixer[21], and some Transformer-based methods, including SwinIR-light [25] and ELAN[61].
@@ IV-B Main Results
Quantitative evaluation. Table II presents quantitative comparisons for the upscaling factors of 2x, 3x, and 4x on five test datasets. Additionally, we also enumerate the parameter counts and FLOPs for each method. Remarkably, our proposed CFSR outperforms existing advanced CNN-based methods in terms of both PSNR and SSIM across all scales and datasets, and significantly bridges the gap between Transformer-based methods[25, 61]. In detail, when compared to DDistill-SR [59], CFSR maintains superior performance across all scales and datasets, while exhibiting approximately half the computational complexity. Furthermore, let us focus on the 3x super-resolution tasks. In comparison to ShuffleMixer [21], CFSR attains significant performance gains, enhancing PSNR by 0.19 and 0.17, and SSIM by 0.0055 and 0.0014 on the Urban100 and Manga109 test datasets, respectively. Importantly, CFSR achieves these results with reduced parameter counts and computational costs, underscoring its efficiency and effectiveness.
We believe that these results corroborate the potential of CFSR as a resource-efficient and performance-oriented model for lightweight image super-resolution tasks.
@@
Qualitative evaluation. We conducted a visual quality comparison of SR results between our proposed CFSR and five representative models, including CARN [14], IMDN [15], RFDN [60], PAN [18], and ShuffleMixer [21]. The ×$\times 4$ SR results are presented in Figure 5. When we take a closer look at the results in the second column, one can find that our CFSR is able to recover the main structures with sharp textures. Moreover, results of samples ’img_058’ and ’img_024’ in Urban100 dataset showcase that CFSR can obtain clearer edges while others fail.
Furthermore, we use LAM[54], which represents the range of attribution pixels, to visualize receptive fields. Visual results are presented in Figure 6, showing that our CFSR can take advantage of a wider range of information than CARN, IMDN, PAN, ShuffleMixer, obtaining large receptive fields effectively. Let us take the ’Tiger’ image in the second row of Figure 6 as the example. One can find that the proposed CFSR can achieve richer textures with the larger receptive field compared to other advanced methods.
Comparison on unknown degradation. Given that the primary objective of image Super-Resolution (SR) is to address complex real-world degradations and generate visually appealing images, we conduct comprehensive evaluation on the RealSR[71] and unknown DIV2K datasets, known for their intricate degradation patterns. For a fair comparison, we retrain and evaluate the ShuffleMixer model and our CFSR on these datasets, respectively. The results, as detailed in Table IV, clearly demonstrate that our CFSR model outperforms the advanced ShuffleMixer in handling complex degradation tasks.
Furthermore, in subjective image quality assessments, the CFSR model exhibits superior performance, and some results are presented in Figure 7. One can find that our CFSR is able to accurately reconstruct finer details and maintain a high level of clarity, even in areas of intricate patterns and textures. We believe these comparisons showcase the practical effectiveness of the CFSR model in real-world applications.
Method Params. FLOPs RealSR DIV2K unknown
ShuffleMixer 411K 28G 29.16/0.8261 29.17/0.8049
CFSR 307K 17.5G 29.25/0.8266 29.39/0.8111
@@ IV-C Ablation Studies
TABLE V: Comparison of FFN with or without the proposed EDC for ×$\times 4$ SR. We report the results on Urban100 and Manga109 datasets.
@@
w/o 307K 17.5G 26.13/0.7875 30.60/0.9098
w 307K 17.5G 26.21/0.7897 30.72/0.9111
Figure 8: Visual comparisons between latent features learned with or without the EDC (correspond to the ablation results in V). The first column is the reference image sampled from Urban100 dataset. (a) Latent features learned by the vanilla DwConv. (b) Feature maps extracted by our EFN with the EDC layer. One can find that EFN can substantially obtains clear and robust high-frequency information in latent feature maps.
@@
Impact of the kernel size. Our in-depth analysis, as presented in Table Table III, explores the implications of diverse kernel sizes on model performance. The results explicitly underscore that model performance escalates with increasing kernel size. Meanwhile, we present corresponding visualization of the activating pixels by LAM[54] in Figure 4, intuitively showcase the superiority of larger kernels in effectively extending the receptive field.
@@
Let us take a closer look at the results in Table III. Despite larger kernel sizes yielding superior performance, this comes with an associated surge in computational demands. Therefore, we limited our investigation to kernel sizes not exceeding 11, owing to both an evident saturation effect and computational constraints. In detail, a progressive enhancement in PSNR/SSIM values of over 0.05dB/0.001 was recorded as the kernel size grew from 5 to 9 across both Urban100 and Manga109 datasets. This consistent gain emphasizes the potency of large kernels in improving the performance. However, the advancement from a 9 to 11 kernel size revealed a more subdued gain of only 0.03/0.0005 in PSNR/SSIM. Given this negligible enhancement and the saturation observed in escalating kernel size benefits, we determined $9$ as the default size for the CFSR. Furthermore, LAM results of different kernel sizes in LK Mixer are presented in Figure 4. One can find that the proposed LK Mixer in CFSR can effective provide large receptive fields.
Impact of EDC. The efficacy of EDC within the Edge-preserving Feed-forward Network (EFN) is evident from the results presented in Table V. The effectiveness of the EDC within the EFN is demonstrated in Table V. Importantly, the employment of re-parameterization ensures that the model with EDC does not incur additional complexity during the inference phase. To provide a more comprehensive understanding, we intensify our analysis by visualizing the latent feature maps, as presented in Figure 8. Fascinatingly, the incorporation of EDC yields a pronounced improvement in the extraction of high-frequency Information within intermediate features. From Figure 8, one can find that there are more edges and textures in latent feature map of our EFN. This observation significantly demonstrate the effectiveness of our enhanced FFN by EDC.

### v2

@@ III Proposed Method
III-A Overall Network Architecture
In Fig. 2, we illustrate the proposed CFSR framework, which encompasses three pivotal stages: shallow feature extraction, deep feature extraction, and the image reconstruction module. The shallow feature extraction phase is designed to distill low-level image features, such as edges, textures, and fine-grained details, from the input image and map them into a latent space. These features are important for preserving the local structure and details of the image. In the deep feature extraction stage, the model extracts higher-level, more compact feature representations. It captures texture and structure information, which is essential for recovering lost details and enhancing the clarity of the image. Utilizing both shallow and deep features, the image reconstruction module is capable of generating high-resolution images. In the following, we will present the details of these three components.
@@ III-A Overall Network Architecture
Deep feature extraction. Then we use a stack of two basic residual blocks (BRB), which contains several ConvFormer layers (CFL), and a $3\times 3$ convolution layer is added at the end of the BRB to aggregate the local features. Specifically, the detailed implementation of BRB and CFL are presented in Fig. 2, respectively. This process of deep feature extraction is formulated as:
@@ III-B ConvFormer Layer
Large kernel mixer. Self-attention is a powerful feature extractor, but its high computational cost makes it impractical for real-time application. Recent studies have demonstrated the effectiveness of employing large kernel convolutions [31]. In Tab. I, a detailed comparison of the computational complexity between multi-head self-attention (MHSA) [20], local window self-attention (LWSA) [21], and large kernel convolution (LK) [31] is exhibited.
Here the $H,W,C$, and $K$ represent the height, width, channel dimension, and kernel (local window) size, respectively. The comparison reveals that, given identical window and kernel sizes, the computational complexity of the LK token mixer is significantly less than that of MHSA. The complexity of LK approximates to half that of LWSA, while simultaneously exhibiting a reduction in parameter count, where $K<<C$. Consequently, large kernel convolutions offer a more resource-efficient choice for the design of lightweight models. In this paper, we propose a simple yet effective feature mixer module as presented in Fig. 2. Here we take a $1\times 1$ convolution followed by a large kernel convolution as the feature mixing gate. Feature extraction of our LK mixer is formulated as follows:
@@
$F_{gate}$ $=\text{DwConv}_{k\times k}(\text{Conv}_{1\times 1}(F)),$
$F_{out}$ $=\text{Conv}_{1\times 1}(V\odot F_{gate}),$
@@
III-C Edge-preserving Feed-forward Network
In the feed-forward network (FFN) in a vanilla Transformer unit, previous studies have enhanced it by integrating depth-wise convolution, thereby improving the local feature ensemble[34, 33]. Considering that SISR, being an ill-posed problem, aims at learning an inversion from LR to HR, where the high-frequency information is crucial to it [58]. To obtain more high-frequency information in latent features, we propose an edge-preserving feed-forward network (EFN), by our Edge-preserving Depth-Wise Convolution (EDC), as illustrated in Fig. 2. This allows a best of both words for a richer high-frequency information while maintaining local feature ensemble. The implementation of the proposed EFN is as follows:
@@ III-C Edge-preserving Feed-forward Network
$F_{2}$ $=\text{GELU}(F_{1}),$
$F_{EDC}$ $=\text{EDC}(F_{2}),$
$F_{3}$ $=\text{Conv}_{1\times 1}(F_{EDC}),$
@@
Detailed implementation of our EDC is presented in Fig. 3. It takes a multi-branch structure, containing a standard depth-wise convolution (DwConv) and three DwConvs with pre-defined gradient kernels. Denote $K_{3\times 3}\in\mathbb{R}^{C\times 1\times 3\times 3}$ and $B_{3\times 3}$ the learnable kernel weights and bias for the vanilla DwConv, where $C$ presents the output channels and 3 is the spatial size. The feature extraction is formulated as:
@@
where * presents the depth-wise convolution operation.
@@
$\scriptsize K_{Lap4}=\left[\begin{array}[]{ccc}0&+1&0\\ +1&-4&+1\\ 0&+1&0\par\end{array}\right],~{}K_{Lap8}=\left[\begin{array}[]{ccc}+1&+1&+1\\ +1&-8&+1\\ +1&+1&+1\end{array}\right],$ (11)
@@
TABLE II: Quantitative comparison with some state-of-the-art SR approaches on five widely used benchmark datasets. Mult-Adds is evaluated on a $1280\times 720$ HR image. Results of ours are in bold.
@@
RFDN[61] $\times$2 534K 95.0G 38.05/0.9606 33.68/0.9184 32.16/0.8994 32.12/0.9278 38.88/0.9773
RLFN [62] 527K 116G 38.07/0.9607 33.72/0.9187 32.22/0.9000 32.33/0.9299 -/-
DIPNet [45] 527K 119G 37.98/0.9605 33.66/0.9192 32.20/0.9002 32.31/0.9302 38.62/0.9770
@@
DLGSANet [63] 566K 128.1G 38.16/0.9611 33.92/0.9202 32.26/0.9007 32.82/0.9343 39.14/0.9777
Omni-SR [64] 772K 172.1G 38.22/0.9613 33.98/0.9210 32.36/0.9020 33.05/0.9363 39.28/0.9784
@@
DLGSANet [63] 572K 56.8G 34.63/0.9288 30.57/0.8459 29.21/0.8083 28.69/0.8630 34.10/0.9480
Omni-SR[64] 780K 78.0G 34.70/0.9294 30.57/0.8469 29.28/0.8094 28.84/0.8656 34.22/0.9487
@@ IV-A Experimental Setup
Comparison Methods We compare the proposed CFSR with state-of-the-art efficient SR approaches, including IMDN [13], LatticeNet[65], LAPAR[14], SMSR[15], ECBSR[17], DRSAN[60], PAN [16], DDistill[66], RFDN [61], RLFN [62], ShuffleMixer[19], DIPNet [45], SPAN [46] and some recent Transformer-based methods, including SwinIR-light [23], DLGSANet [63] and Omni-SR [64].
@@ IV-B Main Results
Quantitative evaluation. Table II presents quantitative comparisons for the upscaling factors of 2x, 3x, and 4x on five test datasets, including parameter counts and FLOPs for each method. Remarkably, our proposed CFSR outperforms existing advanced CNN-based methods in terms of both PSNR and SSIM across all scales and datasets, and significantly bridges the gap between Transformer-based methods[23, 63, 64]. When compared to DDistill-SR [66], CFSR maintains superior performance across all scales and datasets, while exhibiting approximately half the computational complexity. For 3x super-resolution tasks, CFSR attains significant performance gains over ShuffleMixer [19], enhancing PSNR by 0.19 dB and 0.17 dB, and SSIM by 0.0055 and 0.0014 on the Urban100 and Manga109 test datasets, respectively. Notably, CFSR demonstrates competitive performance against recent NTIRE challenge winners, including RLFN [44], DIPNet [45], and SPAN [46]. Across these comparisons, CFSR consistently achieves comparable or superior performance in terms of PSNR and SSIM.
These results underscore CFSR’s potential as a resource-efficient and performance-oriented model for lightweight image super-resolution tasks. By delivering competitive performance with significantly reduced computational demands, CFSR represents a notable advancement in balancing efficiency and effectiveness in the field of image super-resolution.
@@
Qualitative evaluation. We conducted a visual quality comparison of SR results between our proposed CFSR and five representative models, including CARN [12], IMDN [13], RFDN [61], PAN [16], and ShuffleMixer [19]. The ×$\times 4$ SR results are presented in Fig. 5. When we take a closer look at the results in the second column, one can find that our CFSR is able to recover the main structures with sharp textures. Moreover, results of samples ’img_058’ and ’img_024’ in Urban100 dataset showcase that CFSR can obtain clearer edges while others fail.
Furthermore, we use LAM[59], which represents the range of attribution pixels, to visualize receptive fields. Visual results are presented in Fig. 6, showing that our CFSR can take advantage of a wider range of information than CARN, IMDN, PAN, ShuffleMixer, obtaining large receptive fields effectively. Let us take the ’Tiger’ image in the second row of Fig. 6 as the example. One can find that the proposed CFSR can achieve richer textures with the larger receptive field compared to other advanced methods.
Comparison on unknown degradation. Given that the primary objective of image Super-Resolution (SR) is to address complex real-world degradations and generate visually appealing images, we conduct comprehensive evaluation on the RealSR [75] and unknown DIV2K datasets, known for their intricate degradation patterns. For a fair comparison, we retrain and evaluate the ShuffleMixer model and our CFSR on these datasets, respectively. The results, as detailed in Fig. IV, clearly demonstrate that our CFSR outperforms the advanced ShuffleMixer in handling complex degradation tasks.
Furthermore, in subjective image quality assessments, the CFSR model exhibits superior performance, and some results are presented in Fig. 7. One can find that our CFSR is able to accurately reconstruct finer details and maintain a high level of clarity, even in areas of intricate patterns and textures. We believe these comparisons showcase the practical effectiveness of the CFSR model in real-world applications.
@@ IV-C Ablation Studies
Figure 8: Visual comparisons between latent features learned with or without the EDC (correspond to the ablation results in Tab. V). The first column is the reference image sampled from Urban100 dataset. (a) Latent features learned by the vanilla DwConv. (b) Feature maps extracted by our EFN with the EDC layer. One can find that EFN can substantially obtains clear and robust high-frequency information in latent feature maps.
Method Params. FLOPs RealSR DIV2K unknown
ShuffleMixer 411K 28.0G 29.16/0.8261 29.17/0.8049
CFSR 307K 17.5G 29.25/0.8266 29.39/0.8111
TABLE V: Comparison of different re-parameterized convolution configurations in the EFN for ×$\times 4$ SR. We report the results on Urban100 and Manga109 datasets. Configurations: (a) vanilla depth-wise convolution, (b) repConv with one more depth-wise convolution (c) asymmetric convolution from ACNet, and (d) our proposed EDC.
@@
(a) 307K 17.5G 26.13/0.7875 30.60/0.9098
(b) 307K 17.5G 26.14/0.7875 30.62/0.9100
(c) 307K 17.5G 26.14/0.7876 30.62/0.9099
(d) 307K 17.5G 26.21/0.7897 30.72/0.9111
@@
Impact of the kernel size. Our in-depth analysis, as presented in Tab. III, explores the implications of varying kernel sizes on model performance. The results clearly demonstrate that model performance improves as the kernel size increases. To provide an intuitive understanding, we also present visualizations of the activated pixels using LAM [59] in Fig. 4, showcasing the superiority of larger kernels in effectively extending the receptive field.
@@
Examining the results in Tab. III more closely, we observe that while larger kernel sizes yield superior performance, they also result in increased computational demands. Therefore, we limited our investigation to kernel sizes not exceeding 11 due to both computational constraints and an evident saturation effect. Specifically, a progressive enhancement in PSNR/SSIM values of over 0.05 dB/0.001 was recorded as the kernel size increased from 5 to 9 across both the Urban100 and Manga109 datasets. This consistent gain emphasizes the potency of large kernels in improving performance. However, increasing the kernel size from 9 to 11 resulted in a more modest gain of only 0.03 dB/0.0005 in PSNR/SSIM. Given this negligible improvement and the saturation observed in benefits with larger kernel sizes, we selected 9 as the default kernel size for CFSR. Furthermore, the LAM results for different kernel sizes in the LK Mixer are presented in Fig. 4, illustrating that the proposed LK Mixer in CFSR can effectively provide large receptive fields.
Impact of EDC. To thoroughly evaluate the efficacy of our proposed Edge-preserving Depth-wise Convolution (EDC) within the Edge-preserving Feed-forward Network (EFN), we conducted comprehensive experiments comparing various re-parameterized convolution configurations. The results are presented in Tab. V. We examined four configurations: (a) vanilla depth-wise convolution, (b)repConv with one more depthwise convolution (c) asymmetric convolution, and (d) our proposed EDC. It is important to note that the Edge-oriented Convolution Block (ECB) was originally designed for dense convolution, which introduces significantly more parameters when adapted to our model structure. Our proposed EDC demonstrates substantial performance improvements compared to the vanilla depth-wise convolution (a) and re-parameterization modules without edge priors (b) and (c). These results highlight the superiority of our EDC’s ability in enhancing edge features for super-resolution tasks. The effectiveness of the proposed EFN in leveraging edge extraction priors within the FFN structure is particularly noteworthy. By introducing edge-preserving capabilities, we achieve significant performance gains without incurring additional computational overhead during inference. This balance of improved performance and maintained efficiency is crucial for lightweight super-resolution models.
To provide a more comprehensive understanding, we visualize the latent feature maps in Fig. 8. The incorporation of EDC yields a pronounced improvement in the extraction of high-frequency information within intermediate features. Compared to the vanilla depth-wise convolution in the FFN, Figure 8 clearly shows more distinct edges and textures in the latent feature maps of our EFN. This visual evidence further underscores the substantial benefits gained from integrating edge-preserving priors into the FFN through our re-parameterization approach, highlighting the unique contribution of our method in the context of efficient super-resolution.